# CCFC: Bridging Federated Clustering and Contrastive Learning

### v1

@@ 18            Local information aggregation:
4 Experiment
4.1 Experimental Setup
The proposed method is trained using the Adam optimizer [16], implemented in the PyTorch [25], and the code will be made available.
@@ 4.2 Clustering Performance Comparison
We compare CCFC with five state-of-the-art methods, including k-FED [8], FFCM [27], SDA-FC-KM [31], SDA-FC-FCM [31] and PPFC-GAN [30]. To avoid excessive hyperparameter tuning, we use specific hyperparameter settings for each method across various simulated scenarios within the same dataset.
The clustering performance is assessed based on two metrics, NMI [28] and Kappa [18], as presented in Tab. 1 and Tab. 2, respectively. One can see that: 1) Although the two metrics exhibit different ranks in some federated scenarios, the advantages of the proposed method are quite significant in most scenarios. In some scenarios, its performance scores even double those of the best baseline methods. The most conspicuous case is observed on MNIST with the data heterogeneity $p=0$, where it achieves an NMI score surpassing PPFC-GAN by 0.2654 and a Kappa score exceeding 0.3485. 2) Although the synthetic data aided methods (SDA-FC-KM, SDA-FC-FCM and PPFC-GAN) are immune to, or even can benefit from, the data heterogeneity problem, the proposed method demonstrates a higher potential. 3) Representation learning is poised to become the primary catalyst for progress in federated clustering. The most related method to our proposed CCFC is k-fed, with the main distinction being the inclusion of a representation learning process in CCFC. Surprisingly, this learning process results in substantial NMI score improvements of up to 0.4155 and Kappa score enhancements of 0.4593 on the most conspicuous case. In the next subsection, we will present a more in-depth analysis to demonstrate the benefits of representation learning for federated clustering tasks.
@@
${}^{{\ddagger}}$ ${}^{{\ddagger}}$ ${}^{{\dagger}}$ ${}^{{\dagger}}$ SCFC CCFC
@@
k 3 5 7 9 100 1000 7000
Raw pixels 97.05 96.88 96.94 96.59 94.40 87.31 70.16
PPFC-GAN 95.26 95.20 95.25 95.25 92.73 85.37 71.19
SCFC 98.11 98.23 98.25 98.21 97.54 96.30 94.09
CCFC 98.25 98.33 98.26 98.32 97.86 97.09 97.09
4.3 Representation Learning
We claim that the advantages of PPFC-GAN and CCFC tasks stem from the fact that the learned representations are more clustering-friendly. To substantiate this, we qualitatively and quantitatively investigate the latent representations learned by them on MNIST with the data heterogeneity $p=0$. Moreover, to demonstrate the efficacy of the cluster-contrastive learning, we construct an additional baseline method, named sample-contrastive federated clustering (SCFC), achieved by reverting the cluster-contrastive model to SimSiam. This signifies that, throughout the local training process, SCFC abstains from leveraging the semantic structures discovered by clustering.
Qualitative analysis. As shown in Fig. 4, one can observer that: 1) In the original data space, there are 10 discernible modes, yet many samples from different categories are mixed up together, necessitating representation learning. 2) Representation learning through contrastive learning significantly alleviates the mixing problem, and the cluster-contrastive learning yields more compact representations. 3) Interestingly, the latent representations learned by PPFC-GAN seem to exacerbate the mixing problem, implying that the advantages of PPFC-GAN may not be attributed to the acquisition of more clustering-friendly representations. Hence, to thoroughly figure out this matter, further quantitative analysis is necessary.
As shown in Tab. 4, one can find that: 1) The test performance in the latent spaces is comparable to, or in some cases, inferior to that in the original data space for small values of $k$. However, as $k$ values increase, the merits of latent representation learning progressively come to fruition by mitigating the mixing problem. 2) The proposed method demonstrates superior effectiveness and robustness, better preserving the inherent semantic structure of the data.
@@ 4.3 Representation Learning
4.4 Ablation Study
CCFC comprises three key components: cluster-contrastive learning, FedAvg and the model-contrastive regularization term. To validate the effectiveness of each component, we devise four additional baselines within two distinct federated clustering frameworks:
1) Standalone models. In this set of baselines, each client independently performs sample-contrastive learning (denoted as ${}^{{\ddagger}}$) or cluster-contrastive learning (denoted as ${}^{{\ddagger}}$).
2) FedAvg models. In this set of baselines, each client interactively performs sample-contrastive learning (denoted as ${}^{{\dagger}}$) or cluster-contrastive learning (denoted as ${}^{{\dagger}}$) within the FedAvg framework, without employing the model-contrastive regularization term.
Tab. 3 shows that: 1) The complete model exhibits the most favorable clustering performance, implying that the removal of any component can result in performance degradation. 2) The most substantial improvements arise from the collaborative training among clients, while cluster-contrastive learning and the model-contrastive regularization term can yield further performance improvements.
@@ 4.5 Hyperparameter Sensitivity Analysis
Fig. 5 reveals that: 1) The clustering performance remains robust across a wide range of $\lambda$ under a fixed federated heterogeneous scenario. 2) The optimal value of $\lambda$ differs across different federated scenarios. Consequently, the performance of CCFC within Tab. 1 and Tab. 2 is somewhat underestimated, as we used the same $\lambda$ for different federated scenarios of the same dataset to prevent excessive hpyerparameter tuning. 3) The sensitivity of the model to data heterogeneity is more pronounced compared to $\lambda$.
@@ 4.6 Device Failures
Following [31, 30], we simulate different disconnection scenarios by scaling the disconnection rate, which measures the proportion of disconnected clients among all clients. Throughout the training process, only the connected clients are involved. As shown in Fig. 6, while CCFC exhibits greater sensitivity to device failures compared to the baselines, it remains the best performance in most cases.
In summary: 1) The representations acquired through the proposed cluster-contrastive learning model are more clustering-friendly. 2) Benefiting from the more clustering-friendly representations, CCFC demonstrates superior performance. 3) The considered key components within CCFC are all essential for its superiority. 4) The clustering performance remains robust across a wide range of the tradeoff hyperparameter $\lambda$. 5) While CCFC exhibits greater sensitivity to device failures compared to the baselines, it remains the best performance in most cases.
@@ 5 Conclusion
In this work, we first illustrate the necessity of representation learning in federated clustering tasks through a motivational example, and then highlight the limitations of straightforward extensions from centralized models. To handle these, we propose a cluster-contrastive model for learning more clustering-friendly representations, and introduce a simple yet effective federated clustering method by embedding the cluster-contrastive model into the FedAvg framework with the model-contrastive regularization term.
Comprehensive experiments demonstrate the superiority of the cluster-contrastive model in learning clustering-friendly representations, as well as the excellence of the proposed clustering method in terms of clustering performance and handling device failures. We hope our work will draw the community’s focus to the fundamental role of representation learning in federated clustering, and lay a foundation for future advancements within the community.

### v2

@@ 4.2 Clustering Performance Comparison
Federated clustering, as an emerging field, currently lacks a uniform evaluation framework. For example, none of the benchmark datasets used in studies [1, 2, 51] are consistent, making direct comparisons difficult. Thus, considering the availability of baseline codes and ease of implementation, we select k-FED [1], FFCM [2], SDA-FC-KM [3], SDA-FC-FCM [3] and PPFC-GAN [4] as baselines to validate the effectiveness of CCFC. To avoid excessive hyperparameter tuning, we use specific hyperparameter settings for each method across various simulated scenarios within the same dataset.
The clustering performance is assessed based on two metrics, NMI [10] and Kappa [20], as presented in Table 1 and Table 2, respectively. One can see that: 1) Although the two metrics exhibit different ranks in some federated scenarios, the advantages of the proposed method are quite significant in most scenarios. In some scenarios, its performance scores even double those of the best baseline methods. The most conspicuous case is observed on MNIST with the data heterogeneity $p=0$, where it achieves an NMI score surpassing PPFC-GAN by 0.2654 and a Kappa score exceeding 0.3485. 2) CCFC suffers from data heterogeneity, while the synthetic data-aided methods (SDA-FC-KM, SDA-FC-FCM, and PPFC-GAN) are immune to, and can even benefit from, data heterogeneity. Thus, performing contrastive clustering on shared synthetic data appears to offer a solution to mitigate the data heterogeneity problem without sharing private data. However, the increased utility of synthetic data often comes with a heightened risk of deep data leakage from shared synthetic data. In contrast to synthetic data, which may serve as replicas of real data and thus pose a direct privacy risk [16], sharing model parameters offers a significantly greater degree of privacy preservation. Recall that the core distinction between FC and centralized clustering lies in privacy preservation. Hence, we choose to use the FedAvg framework [17] to perform contrastive clustering, rather than SDA-FC [3]. As widely acknowledged, the issue of data heterogeneity remains unresolved and challenging [15, 18, 19]. Thus, this work will emphasize how contrastive learning aids in FC, while the issue of data heterogeneity is left for future research. 3) Representation learning is poised to become the primary catalyst for progress in federated clustering. The most related method to our proposed CCFC is k-fed, with the main distinction being the inclusion of a representation learning process in CCFC. Surprisingly, this learning process results in substantial NMI score improvements of up to 0.4155 and Kappa score enhancements of 0.4593 on the most conspicuous case. In the next subsection, we will provide a deeper analysis to illustrate how representation learning affects federated clustering tasks.
@@
4.3 Effectiveness Analysis of Representation Learning
In centralized clustering, the strength of the representation learning process is its capacity to map raw data features into more clustering-friendly latent representations that exhibit improved intra-cluster compactness and inter-cluster separability [11, 12]. In FC, we will maintain this belief and validate it by qualitatively and quantitatively analyzing the clustering-friendliness of the learned representations in different federated scenarios.
k 3 5 7 9 100 1000 7000
Raw pixels 97.05 96.88 96.94 96.59 94.40 87.31 70.16
PPFC-GAN 95.26 95.20 95.25 95.25 92.73 85.37 71.19
SCFC 98.11 98.23 98.25 98.21 97.54 96.30 94.09
CCFC 98.25 98.33 98.26 98.32 97.86 97.09 97.09
4.3.1 IID scenario
In this context, we explore both qualitatively and quantitatively the latent representations learned by PPFC-GAN and CCFC on the MNIST dataset with data heterogeneity $p=0$. Moreover, to demonstrate the efficacy of the cluster-contrastive learning, we construct an additional baseline method, named sample-contrastive federated clustering (SCFC), achieved by reverting the cluster-contrastive model to SimSiam. This signifies that, throughout the local training process, SCFC abstains from leveraging the semantic structures discovered by clustering.
Qualitative analysis. As shown in Figure 4, one can observer that: 1) In the raw data space, there are 10 discernible modes, yet many samples from different categories are mixed up together, necessitating representation learning. 2) Representation learning through contrastive learning significantly alleviates the mixing problem, and the cluster-contrastive learning yields more compact representations. 3) Interestingly, the latent representations learned by PPFC-GAN seem to exacerbate the mixing problem, implying that the advantages of PPFC-GAN may not be attributed to the acquisition of more clustering-friendly representations. Hence, to thoroughly figure out this matter, further quantitative analysis is necessary.
As shown in Table 3, one can find that: 1) The test performance in the latent spaces is comparable to, or in some cases, inferior to that in the original data space for small values of $k$. However, as $k$ values increase, the merits of latent representation learning progressively come to fruition by mitigating the mixing problem. 2) The proposed method demonstrates superior effectiveness and robustness, better preserving the inherent semantic structure of the data.
Figure 5: t-SNE visualization of learned representations under data heterogeneity in MNIST (best viewed in color). Each color corresponds to a specific category of digits in MNIST, and each grey marker shape denotes a specific local model trained on private data from the corresponding client. A greater $p$ implies increased data heterogeneity, while a higher CH score denotes more clustering-friendly representations that show improved intra-cluster compactness and inter-cluster separability. However, as $p$ increases, the learned representations gradually lose their clustering-friendly traits.
4.3.2 Non-IID scenarios
While the effectiveness of the representation learning process in CCFC has been confirmed in the simplest federated scenario (IID scenario), and resolving data heterogeneity in FC is not the focus of this work, it is essential to analyze how data heterogeneity impairs this process for the future enhancement of the method.
To this end, we first randomly select two classes of digits in MNIST, each containing 100 samples, distributed across 10 clients with varying data heterogeneity $p$. Then, we visualize the latent representations learned by the local models of CCFC, and report their Calinski-Harabasz (CH) scores [52] based on the true labels to measure the corresponding clustering-friendly properties. A higher CH score denotes more clustering-friendly representations that show improved intra-cluster compactness and inter-cluster separability. As shown in Figure 5, one can see that: 1) The learned representations gradually lose their clustering-friendly properties with an increase in data heterogeneity, indicated by decreased CH scores. 2) Heterogeneous data leads to semantic inconsistency in the latent spaces of local models. In the IID scenario ($p=0$), the latent representations of the same digit, learned by different local models, are intermingled, with their spatial distribution unaffected by the specific client. In contrast, under non-IID conditions (e.g., $p=0.5$), the spatial distribution of the learned representations for the same digit exhibits pronounced client-specific characteristics. 3) Heterogeneous data weakens the discriminative ability of contrastive learning. With the increase in data heterogeneity $p$, the expansion of the major class within local data reduces the available discriminative information during comparison, hindering the differentiation between categories and resulting in overlapping cluster representations, i.e., cluster collapse.
The above discussion explains the reasons behind the overwhelming advantage of the proposed method in the IID scenario and how data heterogeneity undermines its effectiveness. Addressing the challenges of semantic inconsistency and cluster collapse induced by data heterogeneity may further unlock the method’s potential, which is left for future work.
@@ 4.3.2 Non-IID scenarios
SCFC CCFC SCFC CCFC SCFC CCFC
@@
4.4 Ablation Study
Recall that CCFC comprises three key components: cluster-contrastive learning, FedAvg and the model-contrastive regularization term. To validate the effectiveness of each component, we devise four additional baselines within two distinct federated clustering frameworks:
1) Standalone models. In this set of baselines, each client independently performs sample-contrastive learning (denoted as SCFC) or cluster-contrastive learning (denoted as CCFC).
2) FedAvg models. In this set of baselines, each client interactively performs sample-contrastive learning (denoted as SCFC) or cluster-contrastive learning (denoted as CCFC) within the FedAvg framework, without employing the model-contrastive regularization term.
Table 4 shows that: 1) The complete model exhibits the most favorable clustering performance, implying that the removal of any component can result in performance degradation. 2) The most substantial improvements arise from the collaborative training among clients, while cluster-contrastive learning and the model-contrastive regularization term can yield further performance improvements.
@@ 4.5 Hyperparameter Sensitivity Analysis
Figure 7 reveals that: 1) The clustering performance remains robust across a wide range of $\lambda$ under a fixed federated heterogeneous scenario. 2) The optimal value of $\lambda$ differs across different federated scenarios. Consequently, the performance of CCFC within Table 1 and Table 2 is somewhat underestimated, as we used the same $\lambda$ for different federated scenarios of the same dataset to prevent excessive hpyerparameter tuning. 3) The sensitivity of the model to data heterogeneity is more pronounced compared to $\lambda$.
@@ 4.6 Device Failures
Following [3, 4], we simulate different disconnection scenarios by scaling the disconnection rate, which measures the proportion of disconnected clients among all clients. Throughout the training process, only the connected clients are involved. As shown in Figure 6, while CCFC exhibits greater sensitivity to device failures compared to the baselines, it remains the best performance in most cases.
In summary: 1) The representations acquired through the proposed cluster-contrastive learning model are more clustering-friendly. 2) Benefiting from the more clustering-friendly representations, CCFC demonstrates superior performance. 3) With the rise of data heterogeneity, the learned representations progressively lose their clustering-friendly traits, as evidenced by semantic inconsistency in the latent spaces of local models and overlapping cluster representations (cluster collapse). 4) The considered key components within CCFC are all essential for its superiority. 5) The clustering performance remains robust across a wide range of the tradeoff hyperparameter $\lambda$. 6) While CCFC exhibits greater sensitivity to device failures compared to the baselines, it remains the best performance in most cases.
@@ 5 Conclusion
As widely acknowledged, the performance of federated methods can be hindered by the issue of data heterogeneity, and this problem remains unresolved and challenging [15, 18, 19]. Thus, this work focuses on how contrastive learning aids in FC and attempts to analyze how data heterogeneity impairs the clustering-friendliness of the learned representations.
To this end, we first propose a cluster-contrastive model for learning more clustering-friendly representations, and introduce a simple yet effective federated clustering method by embedding the cluster-contrastive model into the FedAvg framework. Comprehensive experiments demonstrate the superiority of the cluster-contrastive model in learning clustering-friendly representations, as well as the excellence of the proposed clustering method in terms of clustering performance and handling device failures. The most related method to our CCFC is k-fed [1], with the main distinction being the inclusion of a representation learning process in CCFC. Surprisingly, this learning process results in substantial NMI [10] score improvements of up to 0.4155 and Kappa [20] score enhancements of 0.4593 on the most conspicuous case. Therefore, representation learning is likely to emerge as the key driver of future advancements in FC.
Then, we qualitatively and quantitatively analyze the clustering-friendliness of the learned representations in different federated scenarios. With the rise of data heterogeneity, the learned representations progressively lose their clustering-friendly traits, as evidenced by semantic inconsistency in the latent spaces of local models and overlapping cluster representations (cluster collapse). Resolving these two unfavorable phenomena may further unleash the potential of CCFC, which is left for future work.
We hope our work will draw the community’s focus to the fundamental role of representation learning in FC, and lay a foundation for future advancements within the community.